How do I find the solution to this x^2-16=0
How do I find to the solutions this x^2+16=-64
How do I solve for X x^2+6x+5=0
How do I factor completely 4s^-49
How do I factor completely x^2+5x+6
How do I find the solutions x^2+3x=0
How do I factor completely 3x^2+5x-3
Thank you for your help

Answer by stanbon(75887)

How do I find the solution to this x^2-16=0
(x-4)(x+4) = 0
x = 4 or x = -4
---------------------------------------------------------
How do I find to the solutions this x^2+16=-64
x^2 + 80 = 0
(x + 4isqrt(5))(x + 4isqrt(5)) = 0
x = -4isqrt(5) or x = -4isqrt(5)
---------------------------------------------------------
How do I solve for X x^2+6x+5=0
(x+5)(x+1) = 0
x = -1 or x = -5
---------------------------------------------------------
How do I factor completely 4s^-49
(2s-7)(2s+7) = 0
s = 7/2 or s = -7/2
---------------------------------------------------------
How do I factor completely x^2+5x+6
(x+3)(x+2) = 0
x = -3 or x = -2
---------------------------------------------------------
How do I find the solutions x^2+3x=0
x(x+3) = 0
x = 0 or x = -3
---------------------------------------------------------
How do I factor completely 3x^2+5x-3
x = [-5 +- sqrt(25 -4*3*-3)]/6
x = [-5 +- sqrt(25+36)]/6
x = [-5 +- sqrt(61)]/6
x = [-5 + sqrt(61)]/6 or x = [-5 - sqrt(61)]/6
==================================================
Cheers,
Stan H.